# Computing an Aircraft’s Gliding Range and Minimal Return Altitude in Presence of Obstacles and Wind

### v1

@@ G.2 Glikonal-M
Figure 14 considers the case of a flat elevation profile, i.e., $h_{\text{min}}=0$ everywhere. Panel (a) shows the function $V_{G}$ via its contour lines and a heatmap. The turquoise lines are the feasible re-entry paths. In panels (b,c) the relative and total error, respectively $(V_{G}-V)/V$ and $V_{G}-V$ are plotted. We observe that the error is maximal along the ±$\pm 45$° (with respect to the x axis) directions, while it is zero on the axes. Notice how the radial patterns of the error are different from those of Glikonal-G. While the presence of wind also plays a role, the difference mainly originates from the methods used to solve the HJB equations. In the case of Glikonal-G, the OUM method requires using a triangulated grid, which is built by adding an edge on one of the two diagonals of each square in the grid. Glikonal-M instead uses the FMM, which uses a square grid without any triangulation.
In the second test we consider a flat elevation profile with a barrier of infinite height. The barrier has two openings where it can be passed ($h_{\text{min}}=0$ in correspondence of the openings). The barrier and the openings are shown in white in panels (b,c) of figure 15. In (a) the heatmap of $V_{G}$ and its contour lines are shown. The feasible return paths, in turquoise, can be seen to pass exclusively from the two openings. The relative error and total error are plotted in (b,c).
@@
Figure 15: Test of Glikonal-M on elevation profile with infinite barrier with two narrow openings. In (b,c) the barrier is visible as a white line broken in two points in correspondence of the openings. In the openings and outside the barrier $h_{\text{min}}=0$. The purple dot indicates the position of the airfield. (a) Heatmap and contour lines of $V_{G}$, the solution outputted by Glikonal-M. The turquoise lines are the feasible re-entry trajectories. (b) Relative error $(V_{G}-V)/V$ in percent. (c) Total error $V_{G}-V$.
To show that $V_{G}$ approaches $V$ when the grid spacing approaches zero, we repeat the previous experiment on a grid with $404\times 404$ nodes and grid spacing $0.25$. Figure 16 illustrates the results. Comparing the panels (a) in 16 and 15 we observe no qualitative difference, however the maximum relative error is reduced by about a factor 4, and so is the maximum total error.
Figure 16: Test of Glikonal-M on elevation profile with infinite barrier with two narrow openings. Same elevation profile as in figure 15. Grid spacing: $0.25$, grid size: $404\times 404$. The purple dot indicates the position of the airfield. (a) Heatmap and contour lines of $V_{G}$, the solution outputted by Glikonal-M. The turquoise lines are the feasible re-entry trajectories. (b) Relative error $(V_{G}-V)/V$ in percent. (c) Total error $V_{G}-V$.
The final test, similar to 13 involves an elevation profile with a barrier in the same position of the previous test, but with a height of 60. Figure 17 (a) depicts the function $V_{G}$ and the feasible trajectories. Notice that if a glider is on the opposite side of the barrier, for a certain range of radials from the airfield, it must fly perpendicular to the barrier and will turn towards the airfield only after passing the barrier. Panels (b,c) depict respectively the relative and total error of $V_{G}$.
Figure 17: Test of Glikonal-M on elevation profile with a finite barrier. The barrier has a height of 60 and traverses the plot vertically in the. The x coordinate of the barrier is 50. The purple dot indicates the position of the airfield. (a) Heatmap and contour lines of $V_{G}$, the solution outputted by Glikonal-M. The turquoise lines are the feasible re-entry trajectories. (b) Relative error $(V_{G}-V)/V$ in percent. (c) Total error $V_{G}-V$.
@@ Appendix H Trajectories in MRAP and GRRP
Do $\gamma_{G}$ and $\gamma_{M}$ follow the same path? The answer in general is no. To prove this we will show a counterexample where the two curves are different. Suppose the glide ratio is one and consider a ’staircase’ elevation profile. Looking at the heatmap in figure 18 the elevation is 0 between 0 and 33 on the x axis, it’s 100 between 34 and 66, and it’s 200 from 66 to 100. In this example the solution of both GRRP and MRAP can be computed analytically. The turquoise lines are the the curves $\gamma_{M}$ while the blue lines are the curves $\gamma_{G}$ (reversed in direction). We observe that the blue lines turn towards the airfield immediately after traversing the first ’step’ of the staircase. Instead the turquoise line keep going straight and turn after traversing the second step.
@@
Let us look at GRRP, and consider a point $\bm{y}$ with $y_{1}>66$. The curve $\gamma_{G}$ starting from $\bm{y}$ with altitude $V(\bm{y})=200+(y_{1}-66)$ must initially go straight towards the point $(66,y_{2})$, which will be reached at exactly an altitude of $200$. After that however $\gamma_{G}$ turns towards the airfield, as there are no more obstacles and therefore proceeds in a straight line.
@@
notice that in absence of obstacles the two trajectories coincide. This can be seen also from the analytical solutions presented in B.

### v2

@@ G.2 Glikonal-M
Figure 14 considers the case of a flat elevation profile, i.e., $h_{\text{min}}=0$ everywhere. Panel (a) shows the function $V_{G}$ via its contour lines and a heatmap. The turquoise lines are the feasible re-entry paths. In panels (b,c) the relative and total error, respectively $(V_{G}-V)/V$ and $V_{G}-V$ are plotted. We observe that the error is maximal along the ±$\pm 45$° (with respect to the x-axis) directions, while it is zero on the axes. Notice how the radial patterns of the error are different from those of Glikonal-G. While the presence of wind also plays a role, the difference mainly originates from the methods used to solve the HJB equations. In the case of Glikonal-G, the OUM method requires using a triangulated grid, which is built by adding an edge on one of the two diagonals of each square in the grid. Glikonal-M instead uses the FMM, which uses a square grid without any triangulation.
In the second test, we consider a flat elevation profile with a barrier of infinite height. The barrier has two openings where it can be passed ($h_{\text{min}}=0$ in correspondence of the openings). The barrier and the openings are shown in white in panels (b,c) of figure 15. In (a) the heatmap of $V_{G}$ and its contour lines are shown. The feasible return paths, in turquoise, can be seen to pass exclusively from the two openings. The relative error and total error are plotted in (b,c).
@@
Figure 15: Test of Glikonal-M on elevation profile with an infinite barrier with two narrow openings. In (b,c) the barrier is visible as a white line broken in two points in correspondence of the openings. In the openings and outside the barrier $h_{\text{min}}=0$. The purple dot indicates the position of the airfield. (a) Heatmap and contour lines of $V_{G}$, the solution outputted by Glikonal-M. The turquoise lines are the feasible re-entry trajectories. (b) Relative error $(V_{G}-V)/V$ in percent. (c) Total error $V_{G}-V$.
To show that $V_{G}$ approaches $V$ when the grid spacing approaches zero, we repeat the previous experiment on a grid with $404\times 404$ nodes and grid spacing $0.25$. Figure 16 illustrates the results. Comparing the panels (a) in 16 and 15 we observe no qualitative difference, however, the maximum relative error is reduced by about a factor 4, and so is the maximum total error.
Figure 16: Test of Glikonal-M on elevation profile with an infinite barrier with two narrow openings. Same elevation profile as in figure 15. Grid spacing: $0.25$, grid size: $404\times 404$. The purple dot indicates the position of the airfield. (a) Heatmap and contour lines of $V_{G}$, the solution outputted by Glikonal-M. The turquoise lines are the feasible re-entry trajectories. (b) Relative error $(V_{G}-V)/V$ in percent. (c) Total error $V_{G}-V$.
The final test, similar to 13 involves an elevation profile with a barrier in the same position as the previous test, but with a height of 60. Figure 17 (a) depicts the function $V_{G}$ and the feasible trajectories. Notice that if a glider is on the opposite side of the barrier, for a certain range of radials from the airfield, it must fly perpendicular to the barrier and will turn towards the airfield only after passing the barrier. Panels (b,c) depict respectively the relative and total error of $V_{G}$.
Figure 17: Test of Glikonal-M on elevation profile with a finite barrier. The barrier has a height of 60 and traverses the plot vertically. The x coordinate of the barrier is 50. The purple dot indicates the position of the airfield. (a) Heatmap and contour lines of $V_{G}$, the solution outputted by Glikonal-M. The turquoise lines are the feasible re-entry trajectories. (b) Relative error $(V_{G}-V)/V$ in percent. (c) Total error $V_{G}-V$.
@@ Appendix H Trajectories in MRAP and GRRP
Do $\gamma_{G}$ and $\gamma_{M}$ follow the same path? The answer in general is no. To prove this we will show a counterexample where the two curves are different. Suppose the glide ratio is one and consider a ’staircase’ elevation profile. Looking at the heatmap in figure 18 the elevation is 0 between 0 and 33 on the x-axis, it’s 100 between 34 and 66, and it’s 200 from 66 to 100. In this example, the solution of both GRRP and MRAP can be computed analytically. The turquoise lines are the curves $\gamma_{M}$ while the blue lines are the curves $\gamma_{G}$ (reversed in direction). We observe that the blue lines turn towards the airfield immediately after traversing the first ’step’ of the staircase. Instead, the turquoise lines keep going straight and turn after traversing the second step.
@@
Let us look at GRRP, and consider a point $\bm{y}$ with $y_{1}>66$. The curve $\gamma_{G}$ starting from $\bm{y}$ with altitude $V(\bm{y})=200+(y_{1}-66)$ must initially go straight towards the point $(66,y_{2})$, which will be reached at exactly an altitude of $200$. After that, however, $\gamma_{G}$ turns towards the airfield, as there are no more obstacles, and therefore proceeds in a straight line.
@@
Notice that in the absence of obstacles, the two trajectories coincide. This can be seen also from the analytical solutions presented in B.